# Representation Ensembling for Synergistic Lifelong Learning with Quasilinear Complexity

### v18

@@ 6.2.3 Adversarial experiments
Consider the same CIFAR 10x10 experiments above, but, for Tasks 2 through 9, randomly permute the class labels within each task, rendering each of those tasks adversarial with regard to the first task (because the labels are uninformative). Figure 6A indicates that backward transfer for both SynF and SynN show positive backward transfer even with such label shuffling (the other algorithms did not demonstrate positive backward transfer). Now, consider a Rotated CIFAR experiment, which uses only data from the first task, divided into two equally sized subsets (making two tasks), where the second subset is rotated by different amounts (Figure 6, right). Backward transfer of both SynF and SynN is nearly invariant to rotation angle, whereas the other approaches are far more sensitive to rotation angle. Note that zero rotation angle corresponds to the two tasks having identical distributions. The fact that other algorithms fail to transfer even in this setting suggests that they may not ever be able to positively backwards transfer. See Appendix G.2 for additional experiment using CIFAR 10X10.
Figure 6: Extended CIFAR 10x10 experiments. A. Shuffling class labels within tasks two through nine with 500 samples each demonstrates both SynF and SynN can still achieve positive backward transfer, and that the other algorithms still fail to transfer. B. SynF and SynN are nearly invariant to rotations, whereas other approaches are more sensitive to rotation.
@@ 6.3.1 Spoken Digit
In this experiment, we used the Spoken Digit dataset [57]. As shown in Figure 1 second column, both SynF and SynN show positive backward and forward transfer between the spoken digit tasks, in contrast to other methods, some of which show only forward transfer, others show only backward transfer, with none showing both, and some showing neither. See Appendix G.3 for details of the experiment.
@@ 6.3.2 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of $1000$ food categories from Food2k [58]. FOOD1k 50X20 splits these data into $50$ tasks with $20$ classes each. For each class, we randomly sampled $60$ samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model  Zoo performs the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 1 third column, SynN performs the best among all the algorithms on this large dataset.
@@ 6.3.3 Split Mini-Imagenet
In this experiment, we have used the Mini-Imagenet dataset [54]. The dataset was split into $20$ tasks with $5$ classes each. Each task has $2400$ training samples and $600$ testing samples. As shown in Figure 1 fourth column, we get positive forward and backward transfer for both SynN and SynF. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
@@
In this experiment, we have used 5-dataset [54]. It consists of $5$ tasks from five different datasets: CIFAR-10 [56], MNIST, SVHN [59], notMNIST [60], Fashion-MNIST [61]. All the monochromatic images were converted to RGB format, and then resized to $3\times 32\times 32$. As shown in Appendix Table 4, training samples per task in $5$-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 1 fifth column, SynN and SynF show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SynN and SynF are based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Wyner et al. [62] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1). This is consistent with similar results in single task learning [63, 64, 25]. In lifelong learning, we are often primarily concerned with situations in which we have a small number of samples per task.
@@ 7 Discussion
We introduced quasilinear representation ensembling as an approach to synergistic lifelong learning. Two specific algorithms, SynF and SynN, achieve both forward and backward transfer, by leveraging resources (encoders) learned for other tasks without undue computational burdens. In this paper, we have mainly focused on task-aware setting, because it is simpler. Future work will extend our approach to more challenging task-unaware settings. Ablation experiments with CIFAR 10x10 shows that Forest-based representation ensembling approaches can easily add new resources when appropriate. This work therefore motivates additional research on deep learning to enable dynamically adding resources when appropriate, and resuse the older representations like the moduler methods [65, 66, 26, 49].
@@ Acknowledgements
Appendix A Literature review
Prior work illustrates that ensembling learners can yield huge advantages in a wide range of applications. For example, in classical machine learning, ensembling trees leads to state-of-the-art random forest [38] and gradient boosting tree algorithms [67]. Similarly, ensembling networks shows promising results in various real-world applications [68, 69]. Wang et al. [70] used weighted ensemble of learners in a streaming setting with distribution shift. TrAdaBoost [71] boosts ensemble of learners to enable transfer learning. In continual learning scenarios, many algorithms have been built on these ideas by ensembling dependent representations. For example, Learn++ [72] boosts ensembles of weak learners learned over different data sequences in class incremental lifelong learning settings [73]. Model Zoo [50] uses the same boosting approach in task incremental lifelong learning scenarios.
Another group of algorithms, ProgNN [16] and DF-CNN [17] learn a new “column” of nodes and edges with each new task, and ensembles the columns for inference (such approaches are commonly called ‘modular’ now). The primary difference between ProgNN and DF-CNN is that ProgNN has forward connections to the current column from all the past columns. This creates the possibility of forward transfer while freezing backward transfer. However, the forward connections in ProgNN render it computationally inefficient for a large number of tasks. DF-CNN gets around this problem by learning a common knowledge base and thereby, creating the possibility of backward transfer.
Recently, many other modular approaches have been proposed in the literature that improve on ProgNN’s capacity growth. These methods consider the capacity for each task being composed of modules that can be shared across tasks and grown as necessary. For example, PackNet [66] starts with a fixed capacity network and trains for additional tasks by freeing up portion of the network capacity using iterative pruning. Veniat et al. [26] trains additional modules with each new task, and the old modules are only used selectively. Ostapenko et al. [49] improved the memory efficiency of the modular methods by adding new modules according to the complexity of the new tasks. Mehta et al. [45] proposed non-parametric factorization of the layer weights that promotes sharing of the weights between tasks. However, all of modular methods described above lack backward transfer because the old modules are not updated with the new tasks. Dynamically Expandable Representation (DER) [74] proposed an improvement over the modular approaches where the model capacity is dynamically expanded and the model is fine-tuned by replaying a portion of the old task data along with the new task data. This approach achieves backward transfer between tasks as reported by the authors in the experiments.
Another strategy for building lifelong learning machines is to use total or partial replay [48, 75, 76]. Replay approaches keep the old data and replay them when faced with new tasks to mitigate catastrophic forgetting. However, as we will illustrate, previously proposed replay algorithms do not demonstrate positive backward transfer in our experiments, though they often do not forget as much as other approaches.
Our approach builds directly on previously proposed modular and replay approaches with one key distinction: in our approach, representations are learned independently. Empirically, for low sample sizes random forests (which learn independent trees) typically outperform gradient boosted trees (which learn dependent trees) [64, 77, 78]. Because our approach of representation ensembling is similar to that of random forest, we expect learning independent representations to outperform learning dependent representations in these scenarios as well. This phenomenon is empirically shown in the main text Figure 1. Independent representations also have computational advantages, as doing so merely requires quasilinear time and space, and can be learned in parallel.
Appendix B Evaluation Criteria
@@
Figure 1: Performance summary on vision and audition benchmark datasets using Veniat et al. [26]’s statistics. See Figure 1 for caption details. Note that the results here look nearly identical other than the y-axis labels.
Appendix C Representation Ensembling Algorithms
@@
In this paper, we have proposed two representation ensembling algorithms, Synergistic Forests (SynF) and Synergistic Networks (SynN). The two algorithms differ in their details of how to update encoders and channels, but abstracting a level up they are both special cases of the same procedure. Let SynX refer to any possible synergistic algorithm. Algorithms  123, and  4 provide pseudocode for adding encoders, updating channels, and making predictions for any SynX algorithm. Whenever the learner gets access to a new task data, we use Algorithm 1 to train a new encoder for the corresponding task. We split the data into two portions — one set is used to learn the encoder and the other portion is called the held out or out-of-bag (OOB) data which is returned by Algorithm 1 to be used by Algorithm 2 to learn the channel for the corresponding task. Note that we push the OOB data through the in-task encoder and the whole dataset through the cross-task encoders to update the channel, i.e, learn the posteriors according to the new encoder. Then we use Algorithm 3 to replay the old task data through the new encoder and update their corresponding channels. Finally, while predicting for a test sample, we use Algorithm 4. Given the task identity, we use the corresponding channel to get the average estimated posterior and predict the class label as the $\operatornamewithlimits{argmax}$ of the estimated posteriors.
Algorithm 1 Add a new SynX encoder for a task. OOB = out-of-bag.
@@
7:function SynX.fit($t,(\mathbf{x}^{t},\mathbf{y}^{t})$)
8:     $u_{t},\mathcal{I}_{OOB}^{t}\leftarrow$ encoder.fit($\mathbf{x}^{t}$, $\mathbf{y}^{t}$) $\triangleright$ train an encoder on partitioned data
@@
Algorithm 2 Add a new SynX channel for the current task.
@@
3:(2) $\boldsymbol{u}_{t}=\{u_{t}\}_{t^{\prime}=1}^{t}$triangleright the set of encoders
@@
6:$\boldsymbol{v}_{t}=\{v_{t,t^{\prime}}\}_{t^{\prime}=1}^{t}$triangleright in-task ($t^{\prime}=t$) and cross-task ($t^{\prime}\neq t$) channels for task $t$
7:function SynX.add_channel($t,\boldsymbol{u}_{t},(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$)
8:     $v_{tt}\leftarrow u_{t}$.add_channel($(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$) $\triangleright$ add the in-task channel using OOB data
9:     for $t^{\prime}=1,\ldots,t-1$ do $\triangleright$ update the cross task channels for task $t$
10:         $v_{tt^{\prime}}\leftarrow u_{t^{\prime}}$.add_channel($\mathbf{x}_{t},\mathbf{y}_{t}$)
@@
Algorithm 3 Update SynX channel for the previous tasks.
@@
4:(3) $\mathcal{D}=\{\mathcal{D}^{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright training data for tasks $t^{\prime}=1,\cdots,t-1$
5:$\boldsymbol{v}=\{\boldsymbol{v}_{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright all previous task voters
6:function SynX.update_channel($t,u_{t},\mathcal{D}$)
7:     for $t^{\prime}=1,\ldots,t-1$ do $\triangleright$ update the cross task channels
8:         $v_{t^{\prime}t}\leftarrow u_{t}$.get_channel($\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}}$)
9:     end for
10:     return $\boldsymbol{v}$
11:end function
@@
2:(1) $x\in\mathbb{R}^{p}$triangleright test datum
3:(2) $t$triangleright task identity associated with $x$
4:(3) $\boldsymbol{u}$triangleright all $T$ reperesenters
5:(4) $\boldsymbol{v}_{t}$triangleright channel for task $t$
@@
7:function $\hat{y}=$ SynX.predict($t,x,v_{t}$)
8:     $T\leftarrow$ SynX.get_task_number() $\triangleright$ get the total number of tasks
9:     $\hat{\mathbf{p}}_{t}=\mathbf{0}$triangleright $\hat{\mathbf{p}}_{t}$ is a $K$-dimensional posterior vector
10:     for $t^{\prime}=1,\ldots,T$ do $\triangleright$ aggregate the posteriors calculated from $T$-th task channel
11:         $\hat{\mathbf{p}}_{t}\leftarrow\hat{\mathbf{p}}_{t}+v_{tt^{\prime}}$.predict_proba($u_{t^{\prime}}(x)$)
12:     end for
13:     $\hat{\mathbf{p}}_{t}\leftarrow\hat{\mathbf{p}}_{t}/T$
14:     $\hat{y}=\operatornamewithlimits{argmax}_{i}(\hat{\mathbf{p}}_{t})$triangleright find the index $i$ of the elements in the vector $\hat{\mathbf{p}}_{t}$ with maximum probability
15:     return $\hat{y}$
16:end function
Table 1: Hyperparameters for SynF in CIFAR-10X10 experiments. n_estimators is denoted by $B$, the number of trees, above.
@@
n_estimators ($500$ training samples per task) $10$
n_estimators ($5000$ training samples per task) $40$
max_depth $30$
@@
min_samples_leaf $1$
Table 2: Hyperparameters for SynF in Five Datasets, Split Mini-Imagenet, FOOD1k experiments. n_estimators is denoted by $B$, the number of trees, above. Note that we use the same hyperparameters for all of the aforementioned datasets.
@@
Table 3: Hyperparameters for SynN in CIFAR 10X10, Five Datasets, Split Mini-Imagenet, FOOD1k experiments. Note that we use the same hyperparameters for all of the aforementioned datasets.
Hyperparameters Value
optimizer Adam
learning rate $3\times 10^{-4}$
max_samples (OOB split) $0.67$
K (KNN channel) $\log_{2}(\text{number of samples per task})$
Appendix D Reference Algorithm Implementation Details
The same network architecture was used for all compared deep learning methods. Following van de Ven et al. [48], the ‘base network architecture’ consisted of five convolutional layers followed by two-fully connected layers each containing 2000 nodes with ReLU non-linearities and a softmax output layer. The convolutional layers had $16$, $32$, $64$, $128$ and $254$ channels, they used batch-norm and a ReLU non-linearity, they had a 3x3 kernel, a padding of 1 and a stride of 2 (except the first layer, which had a stride of 1). This architecture was used with a multi-headed output layer (i.e., a different output layer for each task) for all algorithms using a fixed-size network. For ProgNN and DF-CNN the same architecture was used for each column introduced for each new task, and in our SynN this architecture was used for the transformers $u_{t}$ (see above). In these implementations, ProgNN and DF-CNN have the same architecture for each column introduced for each task. Among the reference algorithms, EWC, O-EWC, LwF, SI, Total Replay and Partial Replay results were produced using the repository https://github.com/GMvandeVen/progressive-learning-pytorch. For ProgNN and DF-CNN  we used the code provided in https://github.com/Lifelong-ML/DF-CNN. For all other reference algorithms, we modified the code provided by the authors to match the deep net architecture as mentioned above and used the default hyperparameters provided in the code.
Appendix E Training Time Complexity Analysis

### v19

@@ 6.2.4 Adversarial experiments
Consider the same CIFAR 10x10 experiments above, but, for Task 2 through 9, randomly permute the class labels within each task, rendering each of those tasks adversarial with regard to the first task (because the labels are uninformative). Figure 7A indicates that backward transfer for SiLLy-N show positive backward transfer even with such label shuffling (the other algorithms did not demonstrate positive backward transfer). Now, consider a Rotated CIFAR experiment, which uses only data from the first task, divided into two equally sized subsets (making two tasks), where the second subset is rotated by different amounts (Figure 7, right). Backward transfer of SiLLy-N is nearly invariant to rotation angle, whereas the other approaches are far more sensitive to rotation angle. Note that zero rotation angle corresponds to the two tasks having identical distributions. The fact that other algorithms fail to transfer even in this setting suggests that they may not ever be able to positively backwards transfer. See Appendix A.5 for additional experiment using CIFAR 10X10.
Figure 7: Extended CIFAR 10x10 experiments. A. Shuffling class labels within tasks two through nine with 500 samples each demonstrates SiLLy-N can still achieve positive backward transfer, and that the other algorithms still fail to transfer. B. SiLLy-N is nearly invariant to rotations, whereas other approaches are more sensitive to rotation.
@@ 6.3 Further investigating transfer in additional datasets with more classes, tasks, and/or samples
Figure 8: Performance summary on vision and speech benchmark datasets with varying sample sizes per task. Dataset on the right has more samples per task. Imagenet and five dataset have really high sample size per task which is a non-trivial lifelong learning setting. SiLLy-N performs the best in the low sample size regime on the left two columns which is desirable in lifelong learning. See appendix Figure 3 and 5 for the extended results.
@@ 6.3.1 Spoken Digit
In this experiment, we used the Spoken Digit dataset [55]. As shown in Figure 8 first column, SiLLy-N shows positive backward and forward transfer between the spoken digit tasks, in contrast to other methods, some of which show only forward transfer, others show only backward transfer, with none showing both, and some showing neither. See Appendix A.5 for details of the experiment.
@@ 6.3.2 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of $1000$ food categories from Food2k [56]. FOOD1k 50X20 splits these data into $50$ tasks with $20$ classes each. For each class, we randomly sampled $60$ samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zooperforms the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 8 second column, SiLLy-N performs the best among all the algorithms on this large dataset.
In lifelong learning, we are often primarily concerned with situations in which we have a small number of samples per task. If we have enough samples per tasks, the learning agent need not transfer knowledge from other tasks. However, below we also experiment with non-trivial lifelong learning setting where sample per task is high.
@@ 6.3.3 Split Mini-Imagenet
In this experiment, we have used the Mini-Imagenet dataset [51]. The dataset was split into $20$ tasks with $5$ classes each. Each task has $2400$ training samples and $600$ testing samples. As shown in Figure 1 fourth column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
@@
In this experiment, we have used 5-dataset [51]. It consists of $5$ tasks from five different datasets: CIFAR-10 [53], MNIST, SVHN [57], notMNIST [58], Fashion-MNIST [59]. All the monochromatic images were converted to RGB format, and then resized to $3\times 32\times 32$. As shown in Appendix Table 3, training samples per task in $5$-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 8 fourth column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [60] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1 and 8). This is consistent with similar results in single task learning [61, 62, 25].
6.4 Constant Resource Mode Operation
Figure 9: SiLLy-N in constant resource mode operation. Improvement in Transfer between tasks becomes negligible at nearly $30$ encoders. The user can choose to operate in lower budget with less transfer.
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section 3. We evaluate the performance of SiLLy-N for different number of encoder budget. As shown in Figure 9, performance of SiLLy-N saturates after $30$ encoders, though with only $5$ encoders, still demonstrates forward and backward transfer.
7 SiLLy-F on tabular data
Figure 10: Our proposed approach can be used with random forest as encoders on tabular data (SiLLy-F). SiLLy-F shows more positive forward and backward transfer while operating with less parameters compared to other baseline approaches on tabular data.
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section 3). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having $2000$ nodes, as encoders. As shown in Figure 10, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
@@ 8 Discussion
We introduced representation ensembling as a simple approach for lifelong learning. Two specific algorithms, SiLLy-N and SiLLy-F, achieve both forward and backward transfer, by leveraging resources learned for other tasks without undue computational burdens. In this paper, we have mainly focused on task-aware setting, because it is simpler. Future work will extend our approach to more challenging task-unaware settings. Our code, including code to reproduce the experiments in this manuscript, is available from http://proglearn.neurodata.io/.
@@ Acknowledgements
Appendix A Evaluation Criteria
@@
Figure 1: Performance summary on vision and audition benchmark datasets using Veniat’s [26]’s statistics. See Figure 1 for caption details. Note that the results here look nearly identical other than the y-axis labels.
A.1 Representation Ensembling Algorithms
@@
In this paper, we have proposed two representation ensembling algorithms, Simple Lifelong Learning Networks (SiLLy-N) and Simple Lifelong Learning Forests (SiLLy-F). The two algorithms differ in their details of how to update encoders and channels, but abstracting a level up they are both special cases of the same procedure. Let SiLLy-X refer to any possible representation algorithm. Algorithms  123, and  4 provide pseudocode for adding encoders, updating channels, and making predictions for any SiLLy-X algorithm. Whenever the learner gets access to a new task data, we use Algorithm 1 to train a new encoder for the corresponding task. We split the data into two portions — one set is used to learn the encoder and the other portion is called the held out or out-of-bag (OOB) data which is returned by Algorithm 1 to be used by Algorithm 2 to learn the channel for the corresponding task. Note that we push the OOB data through the in-task encoder and the whole dataset through the cross-task encoders to update the channel, i.e, learn the posteriors according to the new encoder. Then we use Algorithm 3 to replay the old task data through the new encoder and update their corresponding channels. Finally, while predicting for a test sample, we use Algorithm 4. Given the task identity, we use the corresponding channel to get the average estimated posterior and predict the class label as the $\operatornamewithlimits{argmax}$ of the estimated posteriors.
Algorithm 1 Add a new SiLLy-X encoder for a task. OOB = out-of-bag.
@@
7:function SiLLy-X.fit($t,(\mathbf{x}^{t},\mathbf{y}^{t})$)
8:     $u_{t},\mathcal{I}_{OOB}^{t}\leftarrow$ encoder.fit($\mathbf{x}^{t}$, $\mathbf{y}^{t}$) $\triangleright$ train an encoder on training data partitioned into in-bag and OOB samples
@@
Algorithm 2 Add a new SiLLy-X channel for the current task.
@@
3:(2) $\mathcal{U}=\{u_{t^{\prime}}\}_{t^{\prime}=1}^{t}$triangleright the set of encoders
@@
6:$v_{t}$triangleright channel for task $t$
7:function SiLLy-X.add_channel($t,u_{t},(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$)
8:     $v_{t}\leftarrow u_{t}$.add_channel($(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$) $\triangleright$ add the new in-task channel using OOB data
9:     for $t^{\prime}=1,\ldots,t-1$ do
10:         $v_{t}\leftarrow u_{t^{\prime}}$.update_channel($\mathbf{x}_{t},\mathbf{y}_{t},v_{t}$) $\triangleright$ update the channel for task $t$ using the old encoders
@@
Algorithm 3 Update SiLLy-X channel for the previous tasks.
@@
4:(3) $\mathcal{D}=\{\mathcal{D}^{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright training data for old tasks
5:(4) $\mathcal{V}=\{\boldsymbol{v}_{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$triangleright set of all previous task voters
6:$\mathcal{V}=\{\boldsymbol{v}_{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$
7:function SiLLy-X.update_channel($t,u_{t},\mathcal{D},\mathcal{V}$)
8:     for $t^{\prime}=1,\ldots,t-1$ do
9:         $v_{t^{\prime}}\leftarrow u_{t}$.update_channel($\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}},v_{t^{\prime}}$) $\triangleright$ update the old task channels
10:     end for
11:     return $\mathcal{V}$
12:end function
@@
2:(1) $\boldsymbol{x}\in\mathbb{R}^{p}$triangleright test datum
3:(2) $t$triangleright task identity associated with $\boldsymbol{x}$
4:(3) $\mathcal{U}$triangleright set of all $t$ encoders
5:(4) $v_{t}$triangleright channel for task $t$
@@
7:function $\hat{y}=$ SiLLy-X.predict($t,\boldsymbol{x},v_{t}$)
8:     for $t^{\prime}=1,\ldots,t$ do $\triangleright$ get the output $\tilde{\boldsymbol{x}}=\{\tilde{\boldsymbol{x}}_{t^{\prime}}\}_{t^{\prime}=1}^{t}$ from all the encoders
9:         $\tilde{\boldsymbol{x}}_{t^{\prime}}\leftarrow u_{t^{\prime}}.encode(\boldsymbol{x})$
10:     end for
11:     $\hat{\mathbf{p}}\leftarrow v_{t}.predict\_proba(\tilde{\boldsymbol{x}})$triangleright $\hat{\mathbf{p}}$ is a $K_{t}$-dimensional posterior vector
12:     $\hat{y}=\operatornamewithlimits{argmax}(\hat{\mathbf{p}})$triangleright find the index of the elements in the vector $\hat{\mathbf{p}}$ with maximum value
13:     return $\hat{y}$
14:end function
Table 1: Hyperparameters for SiLLy-N in CIFAR 10X10, Five Datasets, Split Mini-Imagenet, FOOD1k experiments. Note that we use the same hyperparameters for all the experiments.
@@
optimizer Adam
learning rate $3\times 10^{-4}$
@@
K (KNN channel) $\log_{2}(\text{number of samples per task})$
Table 2: Hyperparameters for SiLLy-F in tabular CIFAR 10X10 experiments.
@@
A.2 Reference Algorithm Implementation Details
The same network architecture was used for all baseline deep learning methods. Following the work in  [45], the ‘base network architecture’ consisted of five convolutional layers followed by two-fully connected layers each containing 2000 nodes with ReLU non-linearities and a softmax output layer. The convolutional layers had $16$, $32$, $64$, $128$ and $254$ channels, they used batch-norm and a ReLU non-linearity, they had a 3x3 kernel, a padding of 1 and a stride of 2 (except the first layer, which had a stride of 1). This architecture was used with a multi-headed output layer (i.e., a different output layer for each task) for all algorithms using a fixed-size network. For ProgNN and DF-CNN the same architecture was used for each column introduced for each new task, and in our SiLLy-N this architecture was used for the transformers $u_{t}$ (see above). In these implementations, ProgNN and DF-CNN have the same architecture for each column introduced for each task. Among the reference algorithms, EWC, O-EWC, LwF, SI, Total Replay and Partial Replay results were produced using the repository https://github.com/GMvandeVen/progressive-learning-pytorch. For ProgNN and DF-CNN  we used the code provided in https://github.com/Lifelong-ML/DF-CNN. For all other reference algorithms, we modified the code provided by the authors to match the deep net architecture as mentioned above and used the default hyperparameters provided in the code.
A.3 Training Time Complexity Analysis